OpenStudy (anonymous):

9sqrt6 then round three decimals

OpenStudy (anonymous):

got 531441 have 20 problems need help with6

OpenStudy (anonymous):

\(\large 9\sqrt6 \) is already in simplest terms... now use a calculator to round the value to 3 decimal places.

OpenStudy (anonymous):

got531441 if round530441

OpenStudy (anonymous):

those are huge numbers.... the number you should be getting should be less than 25... is this the correct expression: \(\large 9\sqrt6 \) ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

9sqrt6~~

OpenStudy (anonymous):

ok... sqrt6 is about 2.499 so 9 times 2.449 = ????

OpenStudy (anonymous):

22.041

OpenStudy (anonymous):

somewhere there....

OpenStudy (anonymous):

i think they wanted the approximation accurate to 3 decimal places so let your calculator calculate \(\large 9 \times \sqrt6 =\) ???? then round this to 3 decimal places... if you do it this way, you answer would be 22.045

OpenStudy (anonymous):

ok, iam going to calculate

OpenStudy (anonymous):

ok, so if round three decimals it would be22.046

OpenStudy (anonymous):

hmm.. when i let my calculator do it, it gives 22.045407.... so 3 decimal places give 22.045...

OpenStudy (anonymous):

217.474

OpenStudy (anonymous):

thats what, i got the same as u but it said it were wrong and it's217.474

OpenStudy (anonymous):

where did i say it is 217.474 ???

OpenStudy (anonymous):

in mathlab,i know i got the same as you dont know what i am doing wrong because i am using google calculator online

OpenStudy (anonymous):

now i have have to solve9sqrt2

OpenStudy (anonymous):

lol... i used google also...

OpenStudy (anonymous):

i know, i had got the same as u but when i put it end it said it was wrong no offense to uwww.aiunivmathlab.com

OpenStudy (anonymous):

its my hwassignment and every answer i put in it's saying it's wrong, i only got one problem right

OpenStudy (anonymous):

oh wow is iam doing something wrong

OpenStudy (anonymous):

the url does not work.... is it correct? anyway, all the these are calculator prolbems. they want a decimal approximation for the values of the the expressions.... the only way to get it wrong is if you enter the incorrect number of decimal places, or you entered the problem incorrectly...

OpenStudy (anonymous):

is this url correct: www.aiunivmathlab.com ????

OpenStudy (anonymous):

that should work iam on unit5homework

OpenStudy (anonymous):

sorry... i'm getting the login screen...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is this the type of problems on the homework? i don't see how the answer we came up with is incorrect... it is a calculator problem... it could be a problem on their end

OpenStudy (anonymous):

yes, probably

OpenStudy (anonymous):

can we try to do, to go over it together the 24574955 pwaiu.123

OpenStudy (anonymous):

i can help you with a problem but i don't know what "24574955 pwaiu.123" is

OpenStudy (anonymous):

okay 9^2~~

OpenStudy (anonymous):

is that "9 squared" or \(\large 9^2 \) ???

OpenStudy (anonymous):

9^

OpenStudy (anonymous):

says sqrt of 9^2

OpenStudy (anonymous):

the carat symbol "^" is used to refer to exponentiation... so 9^2 means \(\large 9^2 \)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

1458

OpenStudy (anonymous):

ok... so you want "the square root of (9^2) or \(\large \sqrt{9^2} =\) ??? is this correct?

OpenStudy (anonymous):

is the previous post the correct problem?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

yes, but they gave me another one because it were wrong.

OpenStudy (anonymous):

81 is incorrect...

OpenStudy (anonymous):

\(\large \sqrt{9^2}=\sqrt{81}=9 \)

OpenStudy (anonymous):

no it this9sqrt2

OpenStudy (anonymous):

but it asking me to round 3places

OpenStudy (anonymous):

\(\large 9\sqrt2 \approx 12.7279220614 \) if you were to approximate this to 3 decimal places, what should it be?

OpenStudy (anonymous):

12.727911

OpenStudy (anonymous):

3 decimal places means 3 digits (numbers) after the decimal..... so you should have 12.728 you need to look at the 4th digit after the decimal (that would be the "9").. if that digit is 5 or bigger, then add 1 to the third digit. if that digit is less than 5, then the you can disregard all the digits past the third digit

OpenStudy (anonymous):

12.728922

OpenStudy (anonymous):

no... "three decimal digits" means you only have 3 digits to the right of the decimal... so your answer should be 12.728

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

how did you get ~~ on cal

OpenStudy (anonymous):

3sqrt8~~

OpenStudy (anonymous):

i got8.4852814

OpenStudy (anonymous):

round would be8.4952814

OpenStudy (anonymous):

no... 8.48528137424 rounded to 3 decimal places would be 8.485

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

it's still ssaying it's wrong

OpenStudy (anonymous):

it,s not going to let me copy and paste